### v1

# Recommender systems and reinforcement learning for building control and occupant interaction: A text-mining driven review of scientific literature

Abstract
The indoor environment significantly impacts human health and well-being; enhancing health and reducing energy consumption in these settings is a central research focus. With the advancement of Information and Communication Technology (ICT), recommendation systems and reinforcement learning have emerged as promising approaches to induce behavioral changes to improve the indoor environment and energy efficiency of buildings. This study aims to employ text-mining and Natural Language Processing (NLP) techniques to thoroughly examine the connections among these approaches in the context of building control and occupant interaction. The study analyzed around 27,000 articles from the ScienceDirect database, revealing extensive use of recommendation systems and reinforcement learning for space optimization, location recommendations, and personalized control suggestions. Although these systems are broadly applied to specific content, their use in optimizing indoor environments and energy efficiency remains limited. This gap likely arises from the need for interdisciplinary knowledge and extensive sensor data. Traditional recommendation algorithms, including collaborative filtering, content-based and knowledge-based methods, are commonly employed. However, the more complex challenges of optimizing indoor conditions and energy efficiency often depend on sophisticated machine learning (ML) techniques like reinforcement and deep learning. Furthermore, this review underscores the vast potential for expanding recommender systems and reinforcement learning applications in buildings and indoor environments. Fields ripe for innovation include predictive maintenance, building-related product recommendation, and optimization of environments tailored for specific needs, such as sleep and productivity enhancements based on user feedback. The study also notes the limitations of the method in capturing subtle academic nuances. Future improvements could involve integrating and fine-tuning pre-trained language models to better interpret complex texts.
keywords:
Reinforcement Learning , Recommender Systems , Building Operations , Building Energy Efficiency , Word Embeddings , Natural Language Processing
journal: Energy and Buildings
Highlights
Recommender systems and reinforcement learning for building control and occupant interaction: A text-mining driven review of scientific literature
Wenhao Zhang, Matias Quintana, Clayton Miller
  • 27,000 full-text articles related to recommendation systems and the built environment have been extracted for text mining.
  • Detailed relationships among recommender algorithms, system types, input data, interventions, and platforms have been elucidated.
  • Word embeddings model has been used to extract the relationship between keywords.
  • The most commonly used algorithms, well-established applications, and emerging opportunities in building control and occupant interaction have been discussed.
1 Introduction
1.1 Background
The indoor environment significantly affects human health and well-being, with individuals spending on average approximately 86% of their day indoors [1, 2]. Addressing how to improve health and well-being in indoor environments while reducing energy consumption is a central question in contemporary research [3]. In recent years, smart building control has become an important topic in this regard, with extensive studies demonstrating its ability to significantly reduce energy usage while maintaining indoor comfort levels [4, 5, 6, 7]. However, despite these promising results, the adoption rate of smart control technologies remains relatively low [8]. The primary barriers to widespread adoption are the high initial costs and complexity of implementation [9]. Therefore, there remains a need to explore a more feasible, user-friendly, and cost-effective solution.
In the context of the widespread adoption of Artificial Intelligence (AI), the Internet of Things (IoT) devices, and smart mobile devices, a viable approach is the behavior change facilitated by Information and Communication Technology (ICT) [10, 11]. This method leverages recommendation systems and smart devices to influence and change human behavior, thus improving indoor environmental quality and reducing energy consumption [10]. Central to this strategy is sophisticated recommendation algorithms that analyze user preferences and environmental data to provide personalized suggestions. These algorithms have shown success in various sectors, including mobile health, commonly referred to as Just-In-Time Adaptive Interventions (JITAI) in the domain [12], as well as in online shopping [13], entertainment [14, 15], and social networking [16], etc.
In the field of building and built environments, several conventional literature reviews have been conducted recently. These reviews analyze current studies on recommendation systems in various areas, including building performance optimization [17, 18, 19], IoT [20, 21, 22], and smart cities [23, 24, 25, 26]. These systems are typically utilized to optimize indoor environments [27, 28] and enhance building energy efficiency [29, 30, 31, 32]. Predominantly, the methods involve leveraging real-time data and federated learning to train models that generate energy-saving suggestions [33, 34, 35], alongside integrating sensor data with user preference feedback to provide personalized recommendations [36, 37]. Furthermore, reinforcement learning-based recommender systems are widely utilized in this domain, due to their ability to handle multi-objective optimization tasks [38, 39]. For example, multi-agent deep reinforcement learning has been developed to optimize energy consumption, occupant comfort, and air quality simultaneously in commercial buildings [40]. These systems are typically reactive, meaning they respond to user inputs or environmental sensors to provide suggestions. However, proactive recommendation systems have also been explored, which aim to incrementally alter user habits during specific moments based on the theories of micro-moments and habit loops [41]. Additionally, due to the opacity of machine learning (ML) black box models, explainable AI has been used to provide explainable and personalized energy efficiency suggestions [42]. On an urban scale, the application of recommendation systems has also garnered interest, primarily to improve energy efficiency in smart grids [43, 44] and facilitate demand response initiatives [45].
These reviews cover the diverse applications of recommendation systems in various aspects of the built environment. However, due to the limitations of human analytical capacity, conventional literature reviews may not adequately handle the vast volume of publications, struggling to discern the subtle relationships among different studies, particularly in interdisciplinary fields [46, 47]. Typically, these reviews are restricted to analyzing between 80-100 publications. Given the complexity and interdisciplinarity of integrating recommendation systems within the built environment, a more sophisticated approach is essential to efficiently manage and interpret extensive literature datasets. To address these challenges, text mining-based literature review methods emerge as a key tool.
1.2 Text mining-based literature review
Text mining-based literature reviews utilize data-driven technologies to systematically extract and analyze information from a large corpus of text, which is particularly advantageous in fields characterized by rapid technological advancements and extensive academic output [48]. Using natural language processing (NLP) and ML algorithms, text mining can efficiently process text data, including classification, clustering, and association, to identify trends, patterns, and emerging themes that might not be evident through traditional review methods [47, 49].
Previous applications of text mining in literature reviews have demonstrated promising results. Tools such as VOSviewer [50] are commonly used in bibliometric analysis and network visualizations, using paper metadata to create bibliometric networks and density maps between articles in various fields [51, 52, 53]. However, these tools are often constrained by graphical user interfaces (GUIs), which limit the user’s ability to modify or extend beyond the provided algorithms [47]. To address specific text mining needs, researchers have adopted advanced ML techniques such as topic modeling. For example, Latent Dirichlet Allocation (LDA) [54] and the Correlated Topic Model (CTM) [55] have been effectively used to model documents as mixtures of topics [56, 57, 58]. Similarly, [59] employs the BERTopic model [60] to efficiently extract and cluster topics, thus providing deeper insights into textual data. Recent advances in computational power and algorithmic sophistication have also led to the emergence of Large Language Models (LLMs), such as GPT [61] and LLaMA [62], offering an alternative to traditional topic modeling approaches [63]. Furthermore, word embedding technologies such as Word2Vec [64, 65] and GloVe [66] facilitate character-level analysis [47]. These are complemented by tools such as Python and Scikit-Learn [67], RapidMiner Studio [68], N-gram [69], and the Natural Language Toolkit (NLTK) [70] for pre or post-processing, offering a nuanced understanding of word relationships and topic significance [71, 72].
1.3 Review of recommender systems for the built environment using text-mining
This study aims to overcome the challenges and limitations in typical literature reviews while comprehensively examining the interrelationships between recommendation systems in the context of building control and occupant interaction. Specifically, this study utilizes data mining techniques proposed by Abdelrahman et al. [47] to explore the relationships among five distinct categories of keywords. These categories include algorithms, types of recommendation systems, input data, interventions/objectives, and platforms, and Figure 1 illustrates the relational network between these categories. This represents the first effort to apply text mining and NLP to review these elements systematically in the context of the built environment and recommendation systems.
Figure 1: Overview of the categories of concepts analyzed in this text-mining analysis and their relationships with each other.
2 Methodology
The methodology adopted in this study integrates several phases, namely text mining, NLP and visual analytics, which are structured into a systematic six-step process, as illustrated in Figure 2. This comprehensive approach encompasses the following stages: 1) collecting relevant papers via the Elsevier API, 2) extracting and identifying keywords of each category, 3) preprocessing the articles for NLP model training, 4) applying NLP techniques to generate numerical text representations, 5) extracting the relationships among these categories, 6) Creating the relation graph network.
Figure 2: The flowchart shows the methodology used in this research: 1) Identifying the querying keywords of each category, 2) Extracting the relevant articles with their corresponding metadata using Elsevier API, 3) Pre-processing the data, 4) Applying the NLP algorithms, 5) Extracting the relationships among these categories, 6) Similarity matrix visualization.
2.1 Step 1: Article retrieval
ScienceDirect, operated by Elsevier, is a comprehensive academic database and full-text library that serves as major source for scholarly articles [73]. This study uses the Elsevier API [74] to perform full-text searches within the ScienceDirect database. The full text of each article and key metadata, including titles, abstracts, keywords, authors, and publication dates, are extracted and stored in JSON format. This format supports efficient data manipulation and retrieval for subsequent analysis phases.
2.1.1 Search query
The search strategy is designed to extract articles related to recommendation systems within built environments or related to indoor health. The query utilized combines specific terms to refine the focus of the retrieval: (recommendation system OR recommender system OR jitai) AND (energy efficiency OR indoor environment) AND (building OR built environment). The inclusion of terms (buiding OR built environment) ensures that the search results are relevant to the fields of building and built environment studies.
2.1.2 Article filtering
The initial query procedure yielded more than 60,000 articles sourced from more than 1,000 journals (the initial inquiry was conducted in May 2024). A significant proportion of these articles were found to be duplicated and some contained incomplete or erroneous information. Following the removal of these duplicates and erroneous files, the dataset was refined to approximately 27,000 articles. Each article retained includes the full text along with essential metadata such as the author(s), publication year, title, abstract, and journal of publication. Figure 3 displays the distribution of the number of articles per journal and the trends of annual publication. Analysis of these data reveals that most of the articles originate from journals focused on the construction, energy, and environment sectors. In addition, a smaller subset of these articles is derived from medical journals.
Figure 3: Number of collected papers per journal and per year.
2.2 Step 2: Keyword categorization
Upon successful retrieval and download of articles, keywords from each article are extracted and categorized into the five predefined categories outlined in Section 2.2.1, which are essential for subsequent similar word extraction processes. The compound keywords are then transformed into single entities to facilitate the training of NLP models, with constituent words linked by “_”. This modification aims at preserving semantic integrity and will be detailed in the text processing section. For the Named Entity Recognition (NER) task, LLMs have demonstrated commendable accuracy in performing such tasks, as evidenced by existing research [75, 76]. Specifically, this study uses the OpenAI GPT-4 model [77] to carry out NER. A manual review will be conducted following the initial automated classification performed by the LLM.
2.2.1 Definitions
Detailed definitions for each category mentioned in the objectives section are provided below:
  1. Def.1:
    Input data: (input_data) refers to various types of input data employed in recommendation systems. This includes building and environmental data, such as indoor temperature, humidity, and indoor environmental quality; physiological data, such as heart rate, body temperature, and activity patterns; and user context data, such as user profiles and real-time feedback.
  2. Def.2:
    Recommender system: (recommende_system) refers to the types of recommender systems commonly used in the field of the built environment. This encompasses systems like JITAI, context-aware recommender systems for dynamic user environments, and reinforcement learning-based recommender systems that adapt based on user interactions.
  3. Def.3:
    Algorithm: (algorithm) refers to the types of algorithms employed across various recommender systems. Notable examples include NLP, deep learning, and reinforcement learning.
  4. Def.4:
    Objective: (objective) refers to the goals pursued by recommender systems under the context of the built environment. Examples include promoting energy-saving behavior, altering sedentary behaviors, or providing personalized control recommendations.
  5. Def.5:
    Platform: (platform) refers to both the platforms used to collect input data for recommender systems and the platforms on which system outputs are delivered. This encompasses smart wearable devices, as well as traditional computing platforms such as computers and smartphones.
Each of these categories was manually selected from a preliminary survey of the existing literature and extraction of pertinent terms from these texts. At this stage, the extracted articles were ready for preprocessing for the subsequent model training stage.
2.3 Step 3: Text pre-possessing
The pre-processing stage is designed to refine the full-text data for subsequent data mining processes. The preparation involves removing words that could negatively impact model training, merging compound words, and performing word tokenization and lemmatization. Initially, two types of undesirable words are identified: 1) metadata elements such as article or image IDs, author information, hyperlinks, and annotations, including terms such as introduction, methodology, reference, figure, and table, which are repetitive between articles and can introduce bias in subsequent analyses; 2) stop words, which occur frequently but contextually insignificant words such as the, a, in, and of, which if included, could bias the NLP models. NLTK is then used to remove these stop words [70].
Subsequently, the established dictionary of compound words is used to replace compound terms in the articles. This approach allows the model to be trained solely on 1-grams, eliminating the need for additional training of n-gram models to learn these compound terms. An n-gram model refers to a sequence formed by contiguous n words, capable of capturing the relationships between these consecutive words [78]. Typically, training an n-gram model requires preprocessing the text into groups of n words, and for different values of n, the model must be re-trained. For example, to train the NLP model to understand the term building energy model, which constitutes a 3-gram, the input text must be tokenized into groups of three words each for model training. Therefore, after the compound word merging, all n-grams are converted into 1-grams (a single word). This means that each compound word is treated as an independent unit, which simplifies the processing workflow and significantly reduces the time required for model training. Through this method, there is no longer a need to train a dedicated n-gram model to recognize combinations of words. Although this approach only includes words appearing in the keyword section of the papers, it is sufficient to cover the major abbreviations and definitions in the literature, since the keyword section typically contains the core concepts and terms of the articles.
In the final stage, tokenization is performed at the sentence level for each word within the sentences. Since NLP models are case-sensitive, it is necessary to perform lemmatization and stemming to ensure word uniformity. For example, in NLP models, the same word in uppercase and lowercase is treated as distinct entities. To address this, all words are first converted to lowercase. Subsequently, text lemmatization will be employed to extract the common roots of different word forms. After the full text is standardized, it is prepared for the NLP text mining process.
2.4 Step 4: NLP text mining using Word2Vec
Word2Vec is a Python-based word embedding tool that is used to identify semantic similarities between words within a text [64, 65]. This method maps each word to a multidimensional vector, facilitating the measurement of similarity through the cosine similarity of these vectors. Words with vectors that are closer in angle are considered to be more semantically similar.
The objective of training with Word2Vec is to identify an optimal vector representation for each word in the vector space that reflects its semantic meaning. Specifically, Word2Vec utilizes context within sentences, defined as a window that encompasses n words surrounding the target word, to predict the target word or the center word. During training, each word in a sentence is sequentially designated as the center word. The vector representations for both the center word and its surrounding context words are updated based on the discrepancy between the predicted center word vector and its actual vector. Training continues until these discrepancies no longer significantly decrease or model convergence. Convergence implies that the vectors calculated based on surrounding words are sufficiently close to the actual vector of the center word or that further training does not significantly enhance model performance. In this study, a context window of 20 words was used, and each word was assigned a 300-dimensional vector, settings that were determined by hyperparameter tuning.
2.5 Step 5: Extracting the relationship between categories
In the Word2Vec model, the similarity between two words is determined by the frequency of their co-occurrence within similar contexts. For example, if the words thermal_comfort and setpoint appear frequently along with comparable terms, they are represented by vectors that are relatively close in the vector space. The trained word2vec model is used to extract similar words and the similarity between each word in a different category.
An initial step involves the comprehensive extraction of relevant words from each category. Predefined words within each category are used as input to the NLP model to extract words that are semantically similar. This procedure aims to identify and retrieve all relevant words associated with each category from the existing literature. For each word, the model extracts the top 100 similar words. Subsequently, any terms deemed irrelevant are removed using the GPT-4 model. This step ensures that the final dataset is refined and relevant to the specific categories under study. After this preliminary filtering, any remaining words are subjected to manual review and further refinement by human analysts.
Following the extraction of keywords, it is often observed that multiple words may refer to the same concept. In such instances, these words are mapped to a single representative term. For example, the term smart_phone is associated with several synonyms, including android_smartphone, iphone, smartphone, and smart-phone. Variations among these synonyms can involve minor differences, such as the use of underscores (”_”) instead of hyphens (”-”) or the substitution of stage for phase. Another example is the use of acronyms that refer to the same term, such as jitai and just-in-time_adaptive_intervention. Such synonyms and their relationships can be efficiently identified through the similarity matrix in Word2Vec. After compiling a comprehensive list of words and their synonyms from identified categories, the relationships among them are analyzed. The final similarity measure for each list in relation to other words is determined by calculating the average similarity of each word in the list to a target term. This approach ensures a systematic and accurate representation of the terminological relationships within the data set.
Type Description
Content-based system A system that recommends products similar to the user’s past favorites based on feature comparisons [79].
Collaborative filtering-based system A system that utilizes the ratings or behavior of a group of users to make recommendations to similar users [80].
Knowledge-based system A system that uses domain-specific knowledge to provide recommendations [81].
Hybrid recommendation system Combines two or more types of recommender systems to reduce the limitations of any single approach [82].
Social recommendation system Leverages users’ social connections and interactions to recommend items [16].
Context-aware recommendation system Incorporates contexts such as time, location, or environment sensor data to offer more appropriate recommendations [83].
Reinforcement learning-based system A system that dynamically adapts to user preferences over time by continuously learning from user interactions to maximize a reward function [84].
Just-In-Time Adaptive Interventions (JITAI) A system designed to provide intervention on an as-needed basis, adapting in real time to an individual’s changing circumstances and conditions [85].
Table 1: Definition of different types of recommendation systems
3 Results
Table 1 displays the types of recommender systems extracted from the literature and their description. For words extracted in each category, those with a relevance score of less than zero to these recommender system types were eliminated, ensuring that all retained terms positively correlate with the recommender systems.
In this analysis, heat maps are utilized to demonstrate the relevance between pairs of words. Additionally, hierarchical clustering is used to form sub-clusters within each category, which serve to present the compositional structure of each category and enhance the analysis of interrelationships among these clusters. Specifically, Hierarchical Agglomerative Clustering (HAC) utilizing Ward’s method was applied to the word embedding vectors within each category. This approach clusters similar words based on their Euclidean distances in the vector space, facilitating the representation of these clusters in a tree structure known as a dendrogram. Figures 4 and 5 illustrate the data mining results and subclusters within the categories of input data, objectives, algorithms, and platforms.
Figure 4: The hierarchical agglomerative clustering (HAC) of the objective, algorithm, and platform.
Figure 5: The hierarchical agglomerative clustering (HAC) of the input data.
3.1 Input data and objectives of recommender systems for building control and occupant interaction
The initial category extracted the strength of relationships between various input data/objectives and different types of recommender systems. The horizontal axis (Recommender Systems) categorizes the types of recommender systems commonly used in building control and occupant interaction, as identified in the literature. The vertical axis (Input Data/Output Objectives) is organized by categories of terms. The left diagram presents the average relevance of each category of output terms to the recommender systems, while the right diagram provides a detailed classification of outputs.
Figure 6 highlights the significant relevance of real-time and cloud-based data (DT) in this domain, which underscores the sector’s high demand for real-time data, with the majority of recommender system algorithms computed on cloud platforms [10]. The most pertinent data categories for these recommender systems are user context and sensor data (CS). These data types are crucial, as they directly enhance the personalization of recommendations, leveraging real-time environmental and user-specific information to optimize system output. JITAI particularly benefits from high correlations with physiological data (PD) and ecological momentary assessment to provide interventions related to indoor health and behavioral modifications. However, data related to indoor air quality and comfort (IAQ & ITC), such as indoor pollutant concentrations, carbon dioxide levels, ventilation rates, indoor temperatures, noise levels, predicted percentage of dissatisfied occupants, and clothing insulation factor, exhibit lower correlations with recommender systems in the built environment. These data types, despite being crucial for the built environment, are not sufficiently utilized in current recommender systems.
Figure 6: The relationships between input data and different types of recommender systems. The HAC (right) between recommender system type (shown on the X-axis with green highlight) and input data (shown on the Y-axis with blue highlight). The heatmap (left) is a summarized version of the relation map by taking the average of each input data cluster.
Figure 7 provides significant insight into how different types of recommender system align with objectives pertinent to the control of the building and the interaction of the occupants. Regarding the types of recommender systems, it highlights that JITAI is the most widely applicable form of recommendation system within this domain, particularly effective in delivering healthcare and behavioral change interventions (HL). However, reinforcement learning-based recommender systems are not as widely used (showing the lowest relevance to the objectives), but are more closely associated with interventions in indoor environmental comfort and energy performance (IEE). This limited application may be due to the complexity involved in these areas. Regarding the objectives of recommender systems, specific content recommendations (SC), which include suggestions on music, routes, and locations, demonstrate strong associations across all types of recommender systems. However, the analysis reveals a significant gap in the application of recommender systems to objectives directly related to indoor environmental comfort and energy efficiency (IEE). Although these are primary objectives of building control, recommender systems are currently underutilized in these areas.
Figure 7: The relationships between objectives and different types of recommender systems. The HAC (right) between recommender system type (shown on the X-axis with green highlight) and objectives (shown on the Y-axis with red highlight). The heatmap (left) is a summarized version of the relation map by taking the average of each objective cluster.
3.2 Input data for different recommendation objectives
The subsequent comparison provides an overview of the relationship between input data and output in recommender systems, focusing solely on understanding the types of input data required for different outputs. In this analysis, both axes are organized based on the average strength of relationships, with the horizontal axis representing the objectives of the recommender systems (arranged from left to right) and the vertical axis detailing the clusters of input data (arranged from top to bottom).
It can be seen that data are predominantly utilized to enhance indoor comfort and energy performance (IEE), as well as to provide specific content (SC) recommendations. However, data are underused in areas such as health care and lifestyle interventions (HL). On the one hand, some recommender systems frequently utilize data such as temperature settings, work efficiency, daily activities, sedentary behavior, and sleep quality. However, applications that rarely use data include those focused on product recommendations and lifestyle and healthcare interventions.
Figure 8 also reveals diversity in the use of data sources. A key finding is the significant role of ecological momentary assessment data, which is shown to be highly relevant across multiple categories of recommender system objectives. This type of data is crucial as it captures real-time feedback from users, enabling recommendation systems to provide timely and context-specific suggestions that optimize user behavior and system efficiency. Other frequently used data types include smartphone data, mood state, skin temperature, and work productivity. Nonetheless, certain data remain underutilized, including physiological measures (PD) such as heart rate and respiration rate; indoor thermal comfort metrics (ITC) such as clo factor, predicted percentage of dissatisfaction, and metabolic rate; as well as environmental sensor data like air pollutants, weather data, thermal image, and relative humidity.
Figure 8: The relationships between input data (shown on the Y-axis with blue highlight) and objectives (shown on the X-axis with red highlight). Both are grouped based on the HAC and sorted based on the sum of similarities per cluster and per row/column in their clusters.
3.3 Algorithms used in recommender systems
The next comparison uses words related to the algorithm and the types of recommender systems, as shown in Figure 9. This heat map analysis reveals distinct patterns in the applicability of different algorithms across various types of recommender systems for building control and occupant interaction. Predominantly used algorithms include collaborative filtering, content-based, knowledge-driven, fusion-based, and clustering algorithms. NLP exhibits significant effectiveness in content-based and knowledge-driven systems due to its capacity to analyze and interpret complex user data and contextual information. In contrast, regression and optimization algorithms demonstrate a weaker association with most types of recommender systems examined. This may suggest that these more traditional algorithms are less favored for complex or dynamic recommendation tasks that require adaptation to new information or user behaviors. Moreover, algorithms such as reinforcement learning display variable degrees of relationship strength across different types of recommender system, suggesting their applicative versatility and potential underexplored areas for their application.
Figure 9: The relationships between algorithms (shown on the Y-axis with yellow highlight) and different types of recommender systems (shown on the X-axis with green highlight). Both of them are sorted based on the sum of the similarity per each row/column.
3.4 Data collection and delivery platform used in recommender systems
The final visualization in this section focuses on the platforms used in the recommender system. The heat map in Figure 10 illustrates the relationships between the input data types / objectives and the platforms of the recommendation systems. The left side of the figure shows the correlation between the input data and the platforms used to collect these data. Wearable devices and smart home devices exhibit strong correlations with most data types. Head-mounted devices (HMDs), as a new technology, also show a strong correlation with the data. PCs, although capable of collecting building operational data, exhibit weaker connections, possibly due to their limited interaction capabilities and lack of portability.
Figure 10: The heatmap on the left shows the average correlation between each pair of elements from the input data and different platforms. The heatmap on the right shows the average correlation between each pair of elements from the objectives and different platforms.
On the right side, the heatmap illustrates the platforms commonly used to deliver specific content (SC) or behavior interventions (LB). In this domain, smart home devices and wearables remain the primary platforms for interacting with users, likely due to their ubiquitous nature and the continuous data they provide, which are crucial for real-time recommendations and interventions. In contrast, PCs are less utilized for direct interaction in recommendation systems, which can be attributed to their relative immobility and lower user engagement compared to more integrated or personal devices.
4 Discussion
Given the prominence of recommender systems in health management and their growing relevance in the mobile health sector, substantial research has already been published in medical and mobile health journals. Therefore, the subsequent sections will focus on discussing the application of recommender systems in building control and occupant interaction to elucidate their potential and identify gaps in this domain. Specifically, this section will address the following three questions:
  1. 1.
    What are the most commonly used algorithms?
  2. 2.
    What are the most explored applications?
  3. 3.
    What are the emerging application areas in which there are gaps?
Figure 11 presents the relationships between different types of algorithms and the various applications of recommender systems, with a specific emphasis on the application of each algorithm in the field of building control and occupant interaction. Both axes are organized based on the strength of relationships, with the horizontal axis representing the algorithms (arranged from left to right) and the vertical axis detailing the clusters of objectives (arranged from top to bottom). The following subsection will, based on the insights derived from this figure, elucidate key considerations that the research community can address in future studies.
Figure 11: The figure on the left shows the relation between algorithms and recommendation objectives sorted based on usability. The figure on the right shows the relation between algorithms and built environment applications sorted based on the group correlation and element correlation.
4.1 Algorithm selection for applications in building control and occupant interaction
Based on the left side of Figure 11, it can be seen that the algorithms most commonly used for recommender systems include traditional methods such as content-based, collaborative filtering, and knowledge-based systems. However, for applications in building control and occupant interaction, where the variety of data types is extensive, the most commonly used recommendation algorithms are predominantly based on learning algorithms and traditional ML techniques. The five most commonly used recommendation algorithms, illustrated on the right side of Figure 11 (highlighted with black borders), are reinforcement learning, knowledge-based, clustering, fusion-based, and deep learning.
Reinforcement learning-based algorithm operates on the principle of learning optimal actions through interactions with the environment to maximize cumulative rewards [84]. These algorithms continually adjust their strategies based on user interaction feedback, refining their recommendations over time [86]. This adaptive capability makes RL-based algorithms particularly suitable for applications in the built environment, where user preferences, environmental conditions, and energy consumption patterns frequently change [6]. In such applications, value-based methods like Deep Q-Network (DQN) and Double Deep Q-Network (DDQN) algorithms are extensively utilized for energy performance optimization [87] and recommender system design [88]. Owing to their online learning capabilities, which include real-time strategy updates and experience replay mechanisms, these algorithms ensure that the models remain current and efficient [89]. Furthermore, RL-based algorithms are well-suited for addressing multi-objective optimization problems in complex environments. For instance, by integrating XGBoost with DQN, researchers have optimized indoor thermal comfort and energy consumption by predicting the impact of occupants’ behaviors on indoor environments [5]. Despite these advantages, RL algorithms face challenges such as the cold start problem, which is when the system lacks sufficient initial data to make accurate recommendations [38], time-intensive training processes [4], and poor interpretability [89]. Addressing the cold start issue requires substantial data for model training or the incorporation of prior knowledge during the design phase [90, 91]. Moreover, techniques like imitation learning and transfer learning are employed to reduce training times [92, 93], while [94] employs explainable AI to address the challenge of poor interpretability.
Knowledge-based algorithms leverage domain-specific knowledge to generate recommendations, which is crucial when explicit user preferences are unavailable, and decisions must be guided by objective standards or expert knowledge [81]. In the context of the built environment, these algorithms can utilize detailed knowledge about building operations, environmental parameters, and user preferences to recommend optimal strategies [10, 25, 95]. For example, a knowledge-based algorithm can be programmed to maintain acceptable indoor pollutant concentration levels. When sensors detect pollutant levels exceeding predefined thresholds, the system recommends increased ventilation to maintain air quality [20]. Moreover, knowledge-based approaches are often used to enhance other algorithms. For instance, the HeartSteps algorithm, developed to promote physical activity, integrates reinforcement learning with prior knowledge derived from historical data to overcome the cold start problem typical of traditional recommender systems [12].
Clustering algorithms enhance targeted and efficient recommendations by grouping similar items or users based on their attributes or behavioral characteristics using clustering techniques [96]. Unlike collaborative filtering, which relies on user ratings or behavioral similarity [97], clustering algorithms focus on identifying the inherent grouping structure within the data [98, 99], thereby significantly mitigating the cold start problem commonly encountered in collaborative filtering due to data sparsity [100]. For instance, clustering methods like K-means [101] and DBSCAN [102] can perform extensive computations during the preprocessing stage, thereby reducing the need for real-time calculations and improving the operational efficiency of recommender systems. Despite their advantages in handling large-scale datasets, clustering algorithms exhibit limitations in dynamic environments, such as building control or indoor environment optimization [103]. Clustering algorithms typically require major computations, such as determining cluster centers and assigning data points to clusters, to be completed during the preprocessing phase [96, 104]. This means that whenever there are changes in the dataset, such as the addition of building operational data, the clustering model may need to be rerun to reflect these changes. This process is not automatic, contrasting with the real-time responsiveness of collaborative filtering [97].
Fusion-based (or Hybrid) recommendation algorithms offer a solution to address the complexity of data and the diversity of requirements by integrating multiple data sources and recommendation techniques [82], such as collaborative filtering, content-based recommendation, and knowledge-driven approaches. These methods involve aggregating data representations from various sources prior to decision-making or combining recommendations from multiple sub-recommenders to construct an ensemble of recommendations [105, 106]. For instance, [107] describes a decision support system that provides users with early warnings about disease transmission risks through a data aggregation system, which integrates indoor sensor data, mathematical models, and expert knowledge to generate informed recommendations. Similarly, [108] proposed a smart data fusion method, based on the EM3 platform, which integrates sensor data with user habits and feedback to provide personalized energy-saving recommendations at optimal times. Additionally, [109] presents a hybrid approach that combines ML algorithms with building geometry models to provide energy transition recommendations based on the compatibility of renewable energy sources within buildings. This integration allows the system to capitalize on the complementary strengths of diverse data and algorithms, thereby producing stronger, more comprehensive, and accurate recommendations [110]. However, this approach also introduces challenges related to increased computational demands [111] and the complexity of algorithm tuning [112].
Deep learning-based algorithms utilize complex neural network architectures to model user preferences and item characteristics within a recommendation framework [113]. These systems employ multi-layer nonlinear processing units to learn deep representations of features from complex user behaviors, item properties, and textual descriptions, enabling them to discern intricate patterns in large-scale data that traditional algorithms might overlook [114, 115]. For instance, in [116], deep learning models were used to address significant cold start issues in collaborative filtering systems by embedding information about users and their preferences directly into the model. Moreover, deep learning techniques are frequently applied to process unstructured data, such as text, images, and audio [10, 117]. Such as Convolutional Neural Networks (CNNs) [118, 119] are used for image recommendations, while Recurrent Neural Networks (RNNs) [120] handle sequential interaction data. Emerging technologies like attention mechanisms and Transformer models [121] have been widely adopted in deep learning-based recommender systems [114]. These models focus more on the most pertinent parts of the input data, dynamically adjusting the importance weights of different segments to optimize information processing and feature extraction [122]. For instance, in [123], the incorporation of time-aware self-attention mechanisms enhances the precision of capturing user preferences, significantly improving the performance of recommendation models.
4.2 Well-established applications of recommender systems in building control and occupant interaction
From the specific perspective of applications using recommender systems shown on the right side of Figure 11, location-based recommendations, together with the corresponding optimization of space utilization, community recommendations, energy efficiency strategy recommendations, and personalized control recommendations aimed at optimizing indoor environments, are among the most popular (highlighted with red border).
Location recommendation aims to suggest optimal places to users based on their preferences, behaviors, and contextual information [124]. These systems leverage various data sources, including geographic information, user location history, and social networks, to generate personalized location recommendations [125]. For instance, an enhanced recommendation method employs Item-based Collaborative Filtering to recommend new points of interest based on non-spatial attributes similar to previously liked locations [126, 127]. However, traditional collaborative filtering approaches often face the challenge of high sparsity in the user-item matrix, which significantly limits their effectiveness [128]. To address this issue, researchers have proposed methods that integrate various Probability Matrix Factorization models to reduce dependency on specific data types [129]. Additionally, the LCARS methodology proposed in [128] combines offline modeling and online recommendations not only to address the data sparsity issue but also to enhance explainability. In the context of building management, location recommendation systems are also frequently employed for internal space optimization to enhance energy efficiency, comfort, and functionality by optimizing the use of indoor spaces [10]. Given the complexity and diversity of data, methods based on deep learning and reinforcement learning are widely applied in this domain [38, 130]. These algorithms often consider multidimensional factors such as users’ occupancy patterns, environmental conditions, and personal preferences to provide effective solutions for indoor space management and optimization [131].
Personalized control recommendations is another principal application. In the field of built environment, these systems are commonly employed to optimize indoor conditions such as temperature, lighting, and air quality, involving the processing and analysis of time-series data on environmental parameters and user behaviors [10]. Traditional recommendation algorithms, like conventional collaborative filtering, typically lack the capability to handle time-series data and do not possess mechanisms for balancing multiple objectives in complex scenarios [132]. Consequently, personalized control recommendation systems for building energy patterns and indoor environment optimization frequently utilize reinforcement learning, neural networks, and advanced collaborative filtering algorithms to overcome the limitations of conventional recommendation algorithms and address the overload problem [133]. For example, [133] introduced a deep reinforcement learning framework designed to autonomously adjust intelligent lighting systems, enhancing energy efficiency and user satisfaction. Some studies have also replaced collaborative filtering with mathematical models, such as the Manhattan distance, for making recommendations [134]. However, recommender systems face significant challenges in adapting to dynamic environments and balancing multiple objectives [28].
Building energy optimization recommender systems enhance energy efficiency and sustainability by analyzing data from various sources, including occupancy patterns, environmental sensors, and historical energy usage [10]. In the context of global energy challenges, these systems represent a significant future development trajectory. Despite extensive research focused on optimizing building controls, the adoption rates of IoT home appliances and whole-home control systems remain low in residential buildings, leading to limited practical deployment [135]. Consequently, recommender systems that promote behavior change are increasingly viewed as a more effective alternative to conventional smart control systems. According to Figure 11, current methodologies include RL, which dynamically adjusts energy recommendations by learning from interactions with the environment [38]; clustering algorithms, which predict optimal energy settings based on historical data and user behavior [28]; and rule-based models, which offer the advantage of implementing structured decision-making processes that can straightforwardly incorporate expert knowledge and predefined rules into the energy management system [136]. Additionally, optimization algorithms, particularly those integrated with building energy models, are used to address complex multi-objective optimization tasks [137, 138]. The recent focus in this domain includes shifting towards occupant-centered recommendation, integrating personal data to enhance user interactions, designing interaction platforms, and addressing user acceptance and privacy concerns [10].
Community recommendation systems enhance social interaction and community engagement by providing personalized recommendations based on data related to activities, interactions, and preferences [139]. These systems have been widely implemented across various sectors, including community event recommendations, optimization of facility usage, facilitation of resident interactions, and emergency response and safety management [24]. Traditionally, these systems rely on established recommendation algorithms such as collaborative filtering, content-based filtering, and hybrid methods. For instance, [140] proposed cohort comfort models that aims to use occupant similarity to predict an individual’s thermal preference. Furthermore, NLP techniques, such as topic modeling [141], are frequently utilized to enhance the processing of unstructured data, which includes text and conversational data. Additionally, advancements in LLMs, such as GPT (Generative Pre-trained Transformer) APIs, are being leveraged for complex semantic analysis [142]. Ongoing research in this field aims to augment the contextual awareness of these algorithms while addressing challenges related to data privacy and security [143].
4.3 Emerging applications and opportunities
In Figure 11, the right side illustrates that certain recommendation objectives have notably low correlations with the underlying algorithmic concepts (highlighted with pink border). These weak links suggest that these areas may not be adequately addressed by current algorithmic frameworks, highlighting existing gaps and opportunities for innovation. These objectives include building fault detection and predictive maintenance, product recommendations, and specific indoor environment recommendations such as improving sleep quality and work efficiency.
Automated fault detection and diagnostics (AFDD) has rapidly evolved since the early 1990s [47]. Traditionally, fault detection and diagnosis relied on rule-based algorithms and empirical formulas to generate reactive and preventive maintenance plans [144]. However, with recent advances in AI, the focus has shifted towards ML-based systems, which are now recognized for their ability to identify faults and predict maintenance needs in building systems [145]. These systems exhibit exceptional dynamic adjustment and predictive capabilities, enabling timely maintenance and minimizing downtime [144]. By analyzing sensor data and historical maintenance records, these systems can predict potential failures and recommend maintenance actions before issues escalate, marking a significant departure from the traditional methods of periodic checks or reactive maintenance following failures [146]. For instance, integrating ML algorithms with sensor data from HVAC systems allows for real-time insights into equipment health, thereby facilitating proactive maintenance measures [147].
Product recommendation systems are extensively utilized in e-commerce and retail sectors to enhance the user experience by offering personalized product suggestions through algorithms such as collaborative filtering and content-based filtering [148]. However, the potential application scenarios for these systems in the domains of smart buildings and smart homes remain underexplored and underutilized. Current product recommendations typically consider only user behavioral data or basic information such as purchase history, browsing records, or user reviews [149]. Integrating environmental data and user preferences related to energy savings or thermal comfort can significantly improve the relevance and efficiency of recommendations. For instance, when assisting consumers with purchasing appliances, several factors may influence their choices, including the appliance’s energy consumption, manufacturer, price, and its contribution to improving the indoor environment [10]. Additionally, with the deregulation of the retail electricity market, research has also employed recommender systems to provide electricity customers with personalized recommendations for electricity providers and electricity plans tailored to their individual consumption needs [150]. Furthermore, these recommendations can be extended to various aspects of the built environment, such as building design strategies [151] and energy transition solutions [109].
Specific environment optimization through recommender systems represents a future direction for development. Previous discussions have illustrated that recommender systems are utilized to enhance indoor environments by providing personalized control suggestions, primarily focusing on improving indoor comfort and energy efficiency [10]. Recent research has shifted towards offering optimization recommendations for specific settings based on their unique features, such as enhancing sleep quality and work efficiency in sleeping and working environments [152, 153]. These specialized environments require not only environmental data but also an integration of interdisciplinary domain knowledge and individual physiological data or real-time feedback to furnish more accurate recommendations [154]. With the widespread adoption of smart wearable devices and smart home technologies, the real-time acquisition of such data has become feasible. For instance, the Cozie app [155], based on the Apple Watch and Fitbit, can be utilized to gather micro-ecological momentary assessment data and physiological data. By integrating occupant feedback with other sensor data, personalized adjustment recommendations can be created [156]. However, such highly personalized systems pose challenges in the initial data acquisition, processing, and storage of large volumes of sensitive data [10].
5 Conclusion
This study employs text mining techniques to perform a comprehensive review of the literature in the context of building control, occupant interaction, and recommendation systems. By extracting and analyzing trends, patterns, and relationships from approximately 27,000 articles sourced from Elsevier journals, findings that are typically elusive to traditional review methods are revealed. The results indicate that recommendation systems have been widely applied to specific content, such as space optimization, location recommendations, and personalized control suggestions. However, the application of recommendation system methodologies to optimize indoor environments and energy efficiency is not as common. This discrepancy is likely due to the fact that these applications often require the integration of interdisciplinary domain knowledge and extensive sensor data. Moreover, while various data sources related to user context, activities, and real-time feedback are frequently utilized, other critical data sources, including those related to physiological and indoor environmental factors, are underutilized despite their importance for optimizing specific environments or objectives. In addition, traditional recommendation algorithms such as collaborative filtering, content-based, and knowledge-based methods are widely used in this area. Lastly, due to the robust data collection capabilities of wearable devices and their user-friendly interfaces, they are commonly utilized as platforms to implement recommendation systems.
In addition, this work addresses several key questions regarding the application of recommender systems in building control and occupant interaction. Specifically, it examines the most commonly utilized algorithms, the most explored applications, and emerging opportunities in this field. Among these, deep reinforcement learning stands out as the predominant algorithm for building-related applications, primarily due to its ability to handle complex inputs and self-update based on occupant feedback. Although certain applications such as location recommendation, personalized control, and energy optimization of buildings are widely adopted, significant opportunities for innovation remain in emerging areas. These include areas such as predictive maintenance of building systems, recommendations for building design, appliances, and electricity plans, as well as specific environment optimization, such as those designed to improve sleep and work productivity based on occupant feedback.
This review also acknowledges some limitations. The methodology may not fully comprehend the nuances and contextual subtleties of academic texts, potentially overlooking the impact of pioneering research. Furthermore, NLP systems might misinterpret terms with multiple meanings or highly specialized terminology without sufficient contextual data. Future directions include integrating transformer-based pertrained models (such as BERT [157]) and LLM, which could improve the comprehension of complex academic texts. Furthermore, fine-tuning these models on specific academic datasets could significantly improve the precision with which key information is extracted, potentially transforming the landscape of automated literature reviews.
Reproducibility
This analysis can be reproduced using the data and code found in a GitHub repository [^1] https://github.com/buds-lab/recommender-sys-for-buildings-textmining-review.
CRediT authorship contribution statement
Wenhao Zhang: Writing – original draft, Visualization, Software, Methodology, Investigation, Formal analysis, Data curation, Conceptualization. Matias Quintana: Writing – review & editing, Conceptualization. Clayton Miller: Writing – review & editing, Supervision, Resources, Project administration, Funding acquisition, Conceptualization.
Funding
The research was supported by a Ph.D. scholarship from the Singapore Ministry of Education (MOE) and by the National Research Foundation, Prime Minister’s Office, Singapore, under its Campus for Research Excellence and Technological Enterprise (CREATE) program.
Acknowledgements
The authors thank Dr. Mahmoud Abdelrahman for his guidance on using the Elsevier API for article retrieval.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used ChatGPT to perform grammar checking and language polishing. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the published article.